# Characterization of Residual Charge Images in LSST Camera e2v CCDs

### v1

@@ Abstract
LSST Camera CCDs produced by the manufacturer e2v exhibit strong and novel residual charge images when exposed to bright sources. These manifest in images following bright exposures both in the same pixel areas as the bright source, and in the pixels trailing between the source and the serial register. Both of these pose systematic challenges to the Rubin Observatory Legacy Survey of Space and Time instrument signature removal. The latter trail region is especially impactful as it affects a much larger pixel area in a less well defined position. In our study of this effect at UC Davis, we imaged bright spots to characterize these residual charge effects. We find a strong dependence of the residual charge on the parallel clocking scheme, including the relative levels of the clocking voltages, and the timing of gate phase transition during the parallel transfer. Our study points to independent causes of residual charge in the bright spot region and trail region. We propose potential causes in both regions and suggest methodologies for minimizing residual charge. We consider the trade-offs to these methods including decreasing the camera’s full well and dynamic range at the high end. Some of these results and suggestions have been reviewed by the camera commissioning team and may result in changes made to the clocking voltage scheme on the LSST Camera.
@@ 1 INTRODUCTION
The NSF-DOE Vera C. Rubin Observatory Legacy Survey of Space and Time (LSST) depends on the ability to resolve astronomical sources at faint signal levels. When the system is calibrated to image objects as faint as 24.5 magnitude in a 30s exposure over 9.6 square degrees, we will inevitably image bright objects above the dynamic range of the camera. LSST will regularly include bright stellar images and bright transient objects like satellites[1] that exceed the saturation point of the Rubin Observatory LSST Camera (LSSTCam) charge-coupled devices (CCDs). This type of exposure has associated sensor effects that must be characterized and mitigated in the survey for the faint science programs to succeed.
@@
One systematic sensor effect common to CCDs is residual images seen in exposures immediately following a bright or overexposed image. These image “ghosts" are a fairly well documented phenomena. The mechanism through which this occurs is fairly straightforward: When CCDs are exposed to a bright object, not all of the electrons produced are captured in the pixel charge packet or read out in that image. This charge instead persists somewhere in the CCD after the image is read out. On some time scale which is long relative to a single readout, the charge then thermally leaks back into the pixel’s potential well and is read out in subsequent exposures as ghosts of the bright objects [2, 3].
@@
In a more specific example, under some operating conditions it is possible for charge to reach the silicon-silicon oxide (Si-SiO2) interface at the CCD surface where the charge can become trapped. Historically these have been called residual surface images to differentiate them from residual charge images stored in the bulk silicon. However, since the latter is not seen in LSSTCam, we will refer to them simply as residual images or residual charge images. These have also been variously called “image persistence"[4], and the “release of trapped charge"[5] although the mechanism and effect is believed to be the same.
@@
Residual charge images occur when the parallel clocking barrier and collecting voltage levels are configured in such a way that stored charge reaches the surface before it overflows into neighboring pixels in the process called “blooming". The higher collecting voltage level (parallel high), if set high enough, pulls electrons towards the surface. The relative level of the barrier voltage (parallel low) then determines the maximum amount of charge that can be stored in a given pixel before it blooms which we call the “blooming full well". Increasing the number of electrons in a given pixel will also push charge packets further towards the surface. If the full well is large enough that charge reaches the surface, we call this the “surface full well condition". Residual charge can then be addressed proactively with changes in the parallel clocking scheme, or after the fact by raising the substrate voltage to inversion which can provide holes at the surface with which these trapped electrons recombine.
@@
In LSSTCam CCDs there are notable residual charge images in CCDs manufactured by e2v (now Teledyne-e2v)[6, 4]. Residual images have not been observed to the same degree in CCDs manufactured by Imaging Technology Laboratory (ITL) although there are ongoing studies into separate but similar effects due to photoresist impurities. The ITL CCDs are also operated at lower nominal parallel clock levels than e2vs which may account for some of this difference in residual charge[7]. We will show that there are multiple mechanisms that enable residual charge between images in LSSTCam: one that allows charge to be trapped at the same pixels as the imaged bright object, and one that allows charge to be trapped when shifting the charge packets downward toward the serial register during the parallel transfer. These result in distinct manifestations of residual charge images as shown in Figure 1. LSSTCam CCDs all have sixteen separate segments each with their own readout amplifier. The images shown in this study are all images from single amplifier segments, not entire CCDs, as the effects discussed are all confined within a given segment. In addressing these residual charge sources we explore the parameter space of the parallel clocking scheme, how it affects these residual images. We discuss the associated downsides which include excessive leakage current when parallel clock lower voltage is brought too low (less than -6.0 V) and changes to the dynamic range that correspond to changes in the full well level when the parallel clock swing, that is the voltage difference between the high collecting and low barrier voltages, is decreased.
@@
In addition to bright stars, there is also the growing issue of other bright objects which have less well defined locations such as low earth orbit satellites (LEOSats). There are of order 1 million satellites currently filed for launch during the survey[10]. These will start to saturate the CCD sensors starting at 4th magnitude[1]. There are already some satellites recently launched which are as bright as 0th magnitude[11] which will likely bloom the entirety of any of the CCDs which it crosses. Even if only a few tens of thousands of these satellites are bright enough to saturate the full well, we will still see about one of them in every image and they will cross 12 CCD sensors on average. The fact that they cross many pixels or entire sensors exacerbates the issue of residual charge images. The mitigation and removal of residual charge due to LEOSats is then an important issue that demands attention both in the central bright image region and especially in the trail effects that they will cause which will cover large swathes of the camera focal plane.
@@ 2 METHODOLOGY
We developed a series of tests that let us characterize the residual charge effect in all regions. We use an image mask of a single spot that is around 80 pixels in diameter. We take a series of 20 bias images and then expose the sensor to a bright spot source around 4 times the full well saturation with nominal settings in a 15s exposure. The exact level of full well is variable when parameters like voltages or readout clocking are changed. In each test we keep the light source at the same intensity, but depending on the readout settings, the relative full well level changes. In this way we are comparing the response to the same brightness of source object. This will be the best indicator of how the camera will respond to the same on-sky sources for a given change in configurable parameters. We then read out that spot image and immediately follow it with twenty 15 second dark images. We can then look at each subsequent dark to see trends in residual images. We refer to this as our “standard dataset". The sensor at UC Davis has relatively small residual charge compared to some sensors on the main camera. Therefore we repeat this sequence of biases, bright spot image, and dark images 50 times in order to drive up our signal to noise. We keep exposure times at 15s to mimic the cadence of the LSST survey. This means that residual charge images and their scaling to specific parameters will act like those seen in the main survey. It also gives sufficient time resolution to examine the decay of persistence, even if a longer 30s cadence is adopted for LSST.
@@
We also report the median maximum readout we achieved in the coadded bright spot area (labeled ‘1’ in Figure 1) of the oversaturated spot image. This is a stand-in for our full well pixel value which is a measure of our system’s dynamic range which may be impacted by mitigating residual charge. This is not the standard way of calculating the full well value of a sensor which usually entails finding where the photon transfer curve falls off from linear scaling as a function of source brightness. However our method should be a fairly close analog for how the full well value scales in our parameter space and, importantly, it is easily measurable using only our standard dataset. This method is limited by potential variations in gain between different sets of operating parameters.
@@
We can also alter the timing of the CCD readout, and the parallel transfer in particular. In the nominal LSSTCam e2v CCD clocking scheme, when each of the four parallel clocking gates change phase from the low barrier phase to the high collecting phase, there is a short 2660ns period where three of the four gates are held at the collecting voltage at the same time before the next gate changes to the low barrier phase. The gate overlap period accounts for about 23% of the total transfer timing with the nominal readout timing. This standard parallel clocking scheme is shown in Figure 2. This scheme gives the charge ample time to transfer, but may be related to the novel trail region persistence. We do not know the exact mechanism that would cause this but we have a few hypotheses. One possible cause of this would be if, as the three clocks are held high, charge somehow interacts more with the CCD surface. This would cause residual images in the trail region. If this is the case, it is a subtly different mechanism than the standard cause of residual charge images in the spot region during integration. We change the clock timing to decrease the duration of this parallel clock overlap and measure its effect.
@@ 3 RESULTS
Figure 3: Residual charge as a function of time in two frames. Top shows the single image frame which measures the charge in each individual dark following a bright spot exposure. Bottom is the total charge frame which shows the integrated charge seen in the dark images up to that point. Both are useful frames in which to view the exponential thermal decay of electrons from the CCD surface into the pixels. We will mainly use the single image frame in this work unless otherwise noted. This Figure shows the residual charge in the spot region although the trail region residual charge scales similarly. Residual charge has been seen in up to the first ten 15s images and possibly more at a lower level. Changing voltage levels and clocking schemes changes the magnitudes of the charge, but the decay times largely remain unchanged.
We can also examine the structure of the residual images across the affected rows and columns. Taking only the columns where the spot region had been saturated and averaging together the pixels in a given row, we can create a measure of residual charge across each affected row to produce Figure 4. This plot specifically uses data collected with the parallel clock overlap set to around half of its nominal value which just happened to be a dataset where all features and trends in this visualization are accentuated. The same trends exist at other values but are not always as prominent.
@@ 3.1 Parallel Clocking and Residual Charge
Although the dynamic range of LSSTCam will be decreased by lowering the parallel clock swing, the rate at which the maximum pixel readout falls when decreasing the collecting voltage is linear, whereas the residual charge measured decreases exponentially. This is especially pronounced, as evident in Figure 7, in the trail region which is the greater threat to the LSST’s systematic limits. This means that while there is a trade-off to decreasing the parallel high collecting voltage, there is proportionally substantial value gained. Any decrease in the collecting voltage will exponentially decrease the systematics in the survey caused by residual charge images, especially in the trail region. The loss in dynamic range will also be at the bright end. This is important because the main mission of the Rubin observatory is not to get better images of known bright objects, but rather to have a wide survey of dimmer objects and fainter transient objects than have been seen in previous surveys. In presenting this work to the LSST camera team, one of the outcomes was that the camera voltage scheme was changed from the nominal 9.3 V parallel swing to a lower 8.0 V parallel swing by decreasing the high collecting voltage. Other voltages were changed in conjunction to arrive at the configuration called “new" in Table 1. This change resulted in the virtual elimination of residual charge images in the trail region and a large decrease in the spot region. A thorough study still needs to be done using the main LSSTCam sensors to determine the exact quantitative outcomes. We expect the residual charge to scale similarly to that in our system which demonstrated around an 80% decrease in the spot region. This is very likely sensor dependent and our lab only has access to one e2v CCD.
@@ 3.2 Parallel Clock Overlap Time and Residual Charge
Other than changing parallel clocking voltages, the main configurable parameter and independent variable we have to explore is the amount of time that three parallel clocks are held high when gates change phase. We refer to this as the parallel clock overlap. We take our standard dataset with our clocking sequence changed to keep the overall timing to shift down one pixel the same, but changing the overlap by varying amounts from nominal to no overlap. It should be noted that it takes time for the clocks to switch phases and this is not included in the considered overlap time. However this time is on the order of 5-10 ns and should not significantly impact our findings. It also takes a comparable time to switch phase in either direction so the overlap due to this ramp up timing is very limited.
@@
We also operated the CCD in “scan mode"[15, 16], which allows us to read out the raw video pulse for each pixel read into our readout electronics. This revealed that there is no significant change in the signal shape when we change or eliminate the parallel overlap. This indicates that it is not an issue of giving the charge packed adequate time to transfer. It was also considered that the additional residual charge in the trail region is caused by additional time that the charge packet spends in a given region of silicon. This is certainly not the case since the increase in time due to parallel clock overlap is not proportionally long enough.
@@ 4 FUTURE WORK
Residual charge images will appear in the LSST. They are very likely going to be below the detection threshold for a single image and they will not be in the same place for multiple images. When LSSTCam goes on sky it will be necessary to verify that residual charge images are not a significant systematic error. Residuals at the sub-electron level in coadded exposures may be of concern, since surface brightness systematics at this level on several arcsecond scales can affect some science. This will require a fairly in depth study of bright source images and the following images in both the spot and trail regions. This will be confounded by sources in those regions which are not necessarily known or able to be removed from biasing the dataset. The same image reduction algorithm we use in our lab study will not be usable as sky noise and dim sources will offset images from bias frames or dark frames taken before the bright source. That is not an issue in our controlled dark system. This may be mitigated by subtracting the average pixel value of the suspect regions by the average of a similarly sized region of the same image which will have similar sky level and on average the same number of sources. With a sufficiently large sample size, which the LSST will certainly have, we may be able to detect this systematic if it exists and even characterize it for removal in each individual amplifier.

### v2

@@ Abstract
LSST Camera CCDs produced by the manufacturer e2v exhibit strong and novel residual charge images when exposed to bright sources. These manifest in images following bright exposures both in the same pixel areas as the bright source, and in the pixels trailing between the source and the serial register. Both of these pose systematic challenges to the Rubin Observatory Legacy Survey of Space and Time instrument signature removal. The latter trail region is especially impactful as it affects a much larger pixel area in a less well defined position. In our study of this effect at UC Davis, we imaged bright spots to characterize these residual charge effects. We find a strong dependence of the residual charge on the parallel clocking scheme, including the relative levels of the clocking voltages, and the timing of gate phase transition during the parallel transfer. Our study points to independent causes of residual charge in the bright spot region and trail region. We propose potential causes in both regions and suggest methodologies for minimizing residual charge. We consider the trade-offs to these methods including decreasing the camera’s full well and dynamic range at the high end. These results and suggestions have been reviewed by the camera commissioning team and may result in changes made to the clocking voltage scheme on the LSST Camera.
@@ 1 INTRODUCTION
The NSF-DOE Vera C. Rubin Observatory Legacy Survey of Space and Time (LSST) depends on the ability to resolve astronomical sources at faint signal levels[1]. When the system is calibrated to image objects as faint as 24.5 magnitude in a 30s exposure over 9.6 square degrees, we will inevitably image bright objects above the dynamic range of the camera. LSST will regularly include bright stellar images and bright transient objects like satellites[2] that exceed the saturation point of the Rubin Observatory LSST Camera (LSSTCam) charge-coupled devices (CCDs). This type of exposure has associated sensor effects that must be characterized and mitigated in the survey for the faint science programs to succeed.
@@
One systematic sensor effect common to CCDs is residual images seen in exposures immediately following a bright or overexposed image. These image “ghosts" are a fairly well documented phenomena. The mechanism through which this occurs is fairly straightforward: When CCDs are exposed to a bright object, not all of the electrons produced are captured in the pixel charge packet or read out in that image. This charge instead persists within the CCD after the image is read out. On some time scale which is long relative to a single readout, the charge then thermally leaks back into the pixel’s potential well and is read out in subsequent exposures as ghosts of the bright objects [3, 4].
@@
In a more specific example, under some operating conditions it is possible for charge to reach the silicon-silicon oxide (Si-SiO2) interface at the CCD surface where the charge can become trapped. Historically these have been called residual surface images to differentiate them from residual charge images stored in the bulk silicon. However, since the latter is not seen in LSSTCam, we will refer to them simply as residual images or residual charge images. These have also been variously called “image persistence"[5, 6], and the “release of trapped charge"[7] although the mechanism and effect is believed to be the same.
@@
Residual charge images occur when the parallel clocking barrier and collecting voltage levels are configured in such a way that stored charge reaches the surface before it overflows into neighboring pixels in the process called “blooming". The higher collecting voltage level (parallel high), if set high enough, pulls electrons towards the surface. The relative level of the barrier voltage (parallel low) then determines the maximum amount of charge that can be stored in a given pixel before it blooms, which we call the “blooming full well". Increasing the number of electrons in a given pixel will also push charge packets further towards the surface. If the full well is large enough that charge reaches the surface, we call this the “surface full well condition". Residual charge can then be addressed proactively with changes in the parallel clocking scheme, or after the fact by raising the substrate voltage to inversion which can provide holes at the surface with which these trapped electrons recombine.
@@
In LSSTCam CCDs there are notable residual charge images in CCDs manufactured by e2v (now Teledyne-e2v)[5]. Residual images have not been observed to the same degree in CCDs manufactured by Imaging Technology Laboratory (ITL) although there are ongoing studies into separate but similar effects due to photoresist impurities. The ITL CCDs are also operated at lower nominal parallel clock levels than e2vs which may account for some of this difference in residual charge[8]. We will show that there are multiple mechanisms that enable residual charge between images in LSSTCam: one that allows charge to be trapped at the same pixels as the imaged bright object, and one that allows charge to be trapped when shifting the charge packets downward toward the serial register during the parallel transfer. These result in distinct manifestations of residual charge images as shown in Figure 1. LSSTCam CCDs all have sixteen separate segments each with their own readout amplifier. The images shown in this study are all images from single amplifier segments, not entire CCDs, as the effects discussed are all confined within a given segment. In addressing these residual charge sources we explore the parameter space of the parallel clocking scheme, how it affects these residual images. We discuss the associated downsides which include excessive leakage current when parallel clock lower voltage is brought too low (less than -6.0 V) and changes to the dynamic range that correspond to changes in the full well level when the parallel clock swing, that is the voltage difference between the high collecting and low barrier voltages, is decreased.
@@
In addition to bright stars, there is also the growing issue of other bright objects which have less well defined locations such as low earth orbit satellites (LEOSats). There are on the order 1 million satellites currently filed for launch during the survey[11]. These will start to saturate the CCD sensors starting at 4th magnitude[2]. There are already some satellites recently launched which are as bright as 0th magnitude[12] which will likely bloom the entirety of any of the CCDs which it crosses. Even if only a few tens of thousands of these satellites are bright enough to saturate the full well, we will still see about one of them in every image and they will cross 12 CCD sensors on average. The fact that they cross many pixels or entire sensors exacerbates the issue of residual charge images. The mitigation and removal of residual charge due to LEOSats is then an important issue that demands attention both in the central bright image region and especially in the trail effects that they will cause which will cover large swathes of the camera focal plane.
@@ 2 METHODOLOGY
We developed a series of tests that let us characterize the residual charge effect in all regions. We use an image mask of a single spot that is around 80 pixels in diameter. We take a series of 20 bias images and then expose the sensor to a bright spot source around 4 times the full well saturation with nominal settings in a 15s exposure. The exact level of full well is variable when parameters like voltages or readout clocking are changed. In each test we keep the light source at the same intensity, but depending on the readout settings, the relative full well level changes. In this way we are comparing the response to the same brightness of source object. This will be the best indicator of how the camera will respond to the same on-sky sources for a given change in configurable parameters. We then read out that spot image and immediately follow it with twenty 15 second dark images. We can then look at each subsequent dark to see trends in residual images. We refer to this as our “standard dataset". The sensor at UC Davis has relatively small residual charge compared to some sensors on the main camera. Therefore we repeat this sequence of biases, bright spot image, and dark images 50 times in order to drive up our signal to noise. We keep exposure times at 15s to mimic the cadence of the LSST survey. This means that residual charge images and their scaling to specific parameters will act like those seen in the main survey. It also gives sufficient time resolution to examine the decay of residual charge, even if a longer 30s cadence is adopted for LSST.
@@
We also report the median maximum readout we achieved in the coadded bright spot area (labeled ‘1’ in Figure 1) of the oversaturated spot image. This is a stand-in for our full well pixel value which is a measure of our system’s dynamic range which may be impacted by mitigating residual charge. This is not the standard way of calculating the full well value of a sensor which usually entails finding where the photon transfer curve falls off from linear scaling as a function of source brightness. However our method should be a close analog for how the full well value scales in our parameter space and, importantly, it is easily measurable using only our standard dataset. This method is limited by potential variations in gain between different sets of operating parameters.
@@
We can also alter the timing of the CCD readout, and the parallel transfer in particular. In the nominal LSSTCam e2v CCD clocking scheme, when each of the four parallel clocking gates change phase from the low barrier phase to the high collecting phase, there is a short 2660ns period where three of the four gates are held at the collecting voltage at the same time before the next gate changes to the low barrier phase. The gate overlap period accounts for about 23% of the total transfer timing with the nominal readout timing. This standard parallel clocking scheme is shown in Figure 2. This scheme gives the charge ample time to transfer, but may be related to the novel trail region residual charge. We do not know the exact mechanism that would cause this but we have a few hypotheses. One possible cause of this would be that, as the three clocks are held high, charge interacts more with the CCD surface. This would cause residual images in the trail region. If this is the case, it is a subtly different mechanism than the standard cause of residual charge images in the spot region during integration. We change the clock timing to decrease the duration of this parallel clock overlap and measure its effect.
@@ 3 RESULTS
Figure 3: Residual charge as a function of time in two frames. Top shows the single image frame which measures the charge in each individual dark following a bright spot exposure. Bottom is the total charge frame which shows the integrated charge seen in the dark images up to that point. Both are useful frames in which to view the exponential thermal decay of electrons from the CCD surface into the pixels. We will mainly use the single image frame in this work unless otherwise noted. This figure shows the residual charge in the spot region. The trail region residual charge scales similarly. Residual charge has been seen in up to the first ten 15s images and possibly more at a lower level. Changing voltage levels and clocking schemes changes the magnitudes of the charge, but the decay times largely remain unchanged.
We can also examine the structure of the residual images across the affected rows and columns. Taking only the columns the spot region had been saturated and averaging together the pixels in a given row, we can create a measure of residual charge across each affected row to produce Figure 4. This plot specifically uses data collected with the parallel clock overlap set to around half of its nominal value which just happened to be a dataset where all features and trends in this visualization are accentuated. The same trends exist at other values but are not always as prominent.
@@ 3.1 Parallel Clocking and Residual Charge
Although the dynamic range of LSSTCam will be decreased by lowering the parallel clock swing, the rate at which the maximum pixel readout falls when decreasing the collecting voltage is linear, whereas the residual charge measured decreases exponentially. This is especially pronounced, as evident in Figure 7, in the trail region which is the greater threat to the LSST’s systematic limits. This means that while there is a trade-off to decreasing the parallel high collecting voltage, there is proportionally substantial value gained. Any decrease in the collecting voltage will exponentially decrease the systematics in the survey caused by residual charge images, especially in the trail region. The loss in dynamic range will also be at the bright end. This is important because the main mission of the Rubin observatory is not to get better images of known bright objects, but rather to have a wide survey of dimmer objects and fainter transient objects than have been seen in previous surveys. In presenting this work to the LSST camera team, one of the outcomes was that the camera voltage scheme was changed from the nominal 9.3 V parallel swing to a lower 8.0 V parallel swing by decreasing the high collecting voltage. Other voltages were changed in conjunction to arrive at the configuration called “new" in Table 1. This change resulted in the virtual elimination of residual charge images in the trail region and a large decrease in the spot region. A thorough study still needs to be done using the main LSSTCam sensors to determine the exact quantitative outcomes. We expect the residual charge to scale similarly to that in our system which demonstrated around an 80% decrease in the spot region. This is likely sensor dependent and our lab only has access to one e2v CCD.
@@ 3.2 Parallel Clock Overlap Time and Residual Charge
Other than changing parallel clocking voltages, the main configurable parameter and independent variable we have to explore is the amount of time that three parallel clocks are held high when gates change phase. We refer to this as the parallel clock overlap. We take our standard dataset with our clocking sequence changed to keep the overall timing to shift down one pixel the same. We change the overlap by varying amounts from nominal to no overlap. It should be noted that it takes time for the clocks to switch phases and this is not included in the considered overlap time. However this time is on the order of 5-10 ns and should not significantly impact our findings. It also takes a comparable time to switch phase in either direction, so the overlap due to this ramp up timing is very limited.
@@
We also operated the CCD in “scan mode"[16, 17], which allows us to read out the raw video pulse for each pixel read into our readout electronics. We found that there is no significant change in the signal shape when we change or eliminate the parallel overlap. This indicates that it is not an issue of giving the charge packed adequate time to transfer. It was also considered that the additional residual charge in the trail region is caused by additional time that the charge packet spends in a given region of silicon. This is certainly not the case since the increase in time due to parallel clock overlap is not proportionally long enough.
@@ 4 FUTURE WORK
Residual charge images will appear in the LSST. They are very likely going to be below the detection threshold for a single image and they will not be in the same place for multiple images. When LSSTCam goes on-sky it will be necessary to verify that residual charge images are not a significant systematic error. Residuals at the sub-electron level in coadded exposures may be of concern, since surface brightness systematics at this level on several arcsecond scales can affect some science. This will require a fairly in depth study of bright source images and the following images in both the spot and trail regions. This will be confounded by sources in those regions which are not necessarily known or able to be removed from biasing the dataset. The same image reduction algorithm we use in our lab study will not be usable as sky noise and dim sources will offset images from bias frames or dark frames taken before the bright source. That is not an issue in the controlled dark system at UC Davis. This sky offset may be mitigated by subtracting the average pixel value of the suspect regions by the average of a similarly sized region of the same image, which will have a similar sky level and on average the same number of sources. With a sufficiently large sample size, which the LSST will certainly have, we may be able to detect this systematic if it exists and even characterize it for removal in each individual amplifier. If necessary, the process of removal will pose a problem of characterization and subtraction while limiting residual errors. This has been dealt with in other surveys such as Pan-STARRS’ burntool program which removes residual charge signatures using a one-dimensional exponential model[6].